OpenStudy (anonymous):

What are the rules for solving for x? 8^x = (1/32)^(x-2)

OpenStudy (anonymous):

\[8^x = (\frac{ 1 }{ 32 })^{x-2}\]

OpenStudy (fibonaccichick666):

yuck ok so, first question, what can you do to bring the exponents down?

OpenStudy (anonymous):

You have two choices, factor or use logarithms.

OpenStudy (fibonaccichick666):

(I personally prefer logs, and it is not nearly as yucky as it was upon first inspection)

OpenStudy (anonymous):

Factoring would lead to the realization that \(8=2^3\) and \(1/32 = 2^{-5}\).

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ 8^x=(\frac{1}{32})^{x-2} }\) \(\LARGE\color{black}{ 8^x=(\frac{1}{2^5})^{x-2} }\) \(\LARGE\color{black}{ 8^x=(2^{-5})^{x-2} }\) \(\LARGE\color{black}{ (2^3)^x=(2^{-5})^{x-2} }\) \(\LARGE\color{black}{ (2)^{3x}=(2^{-5})^{x-2} }\) \(\LARGE\color{black}{ (2)^{3x}=(2^{})^{-5(x-2)} }\)

OpenStudy (solomonzelman):

hope you can take it from here.

OpenStudy (anonymous):

Yep Solomon Thanks!

OpenStudy (fibonaccichick666):

I don't usually notice that type of stuff, nice catch wio, I need a more accross the board method

OpenStudy (anonymous):

When working with logs, you generally want to factor anyway.

OpenStudy (fibonaccichick666):

sort of, instead of having to notice the ways you can rewrite two, you can take a log of both sides, then by properties of logs, you can bring down the exponents and solve for x

OpenStudy (anonymous):

Ohh ok so Solomon's method would be my number one choice?

OpenStudy (anonymous):

@OwdinMW Maybe you should do your own homework. Maybe you'd learn something that way.

OpenStudy (fibonaccichick666):

it's up to personal preference essentially

OpenStudy (anonymous):

\[\frac{1}{a^m} = a^{-m} \\ (a^m)^n = a^{m \times n} \\ If \ \ \ a^m = a^n, \implies m = n\]

OpenStudy (anonymous):

Alright :) thank you!

OpenStudy (anonymous):

Have you studied logarithms @OdinMW ?

OpenStudy (anonymous):

Yes. I'm just getting into them.

OpenStudy (fibonaccichick666):

If you are just getting into them, chances are you should do the problem both ways to ensure you recieve credit

OpenStudy (anonymous):

\[\log_{b}\frac{ m }{ n } = \log_{b} m - \log_{b} n\]

OpenStudy (fibonaccichick666):

have you learned about how to bring down exponents with logs?

OpenStudy (anonymous):

@FibonacciChick666 yes

OpenStudy (anonymous):

First thing you want to do here is change the fraction into a negative exponent. \[ 8^x=32^{2-x} \]

OpenStudy (fibonaccichick666):

i concur

OpenStudy (anonymous):

Then you would use logs.

OpenStudy (fibonaccichick666):

then what would happen if you took log base 8 of both sides?

OpenStudy (anonymous):

Yes @wio thats something I should have thought of.

OpenStudy (anonymous):

You can keep the base to be whatever you want for the beginning.

OpenStudy (fibonaccichick666):

or ln, either works

OpenStudy (anonymous):

So the log of both sides gives: \[ x\log 8=(2-x)\log32 \]

OpenStudy (fibonaccichick666):

personally, I prefer base 8 here because it makes less work later

OpenStudy (fibonaccichick666):

and I think you can manage from there odin

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Whoops, I wrote down the wrong thing.

OpenStudy (fibonaccichick666):

i got, more along the lines of x=2log32-xlog32

OpenStudy (fibonaccichick666):

the log in this case was base 8

OpenStudy (anonymous):

Im thinking its getting a little too much haha. I would like to know how to solve it with logs but I think Ill just use Solomon's method...

OpenStudy (anonymous):

loooooossssseeeeerrrrrrrrrsssssssssss

OpenStudy (anonymous):

Hmmm:\[ x\log 8=(2-x)\log32\\ x\log8=2\log32-x\log32\\ x\log8+x\log 32 = 2\log 32\\ x(\log8+\log32)=2\log32\\ x = \frac{2\log 32}{\log8+\log32} \]

OpenStudy (fibonaccichick666):

x=2log32-xlog32< this is just a two step problem from here, it's not that bad, I swear :)

OpenStudy (fibonaccichick666):

or what wio did

OpenStudy (anonymous):

Change it to base \(2\), and it becomes: \[ x = \frac{2(5)}{(3)+(5)} \]

OpenStudy (anonymous):

If you had \[ \log_8(32) \]You can always convert this to another base \(c\) using the property: \[ \log_ab = \frac{\log_cb}{\log_ca} \]

OpenStudy (anonymous):

In general, you either want your base to be a prime number, or you just want to use the natural log \(\ln\).

OpenStudy (anonymous):

\[x=32\] basically...right?